OpenStudy (shamil98):

Find all possible rational zeros for the following polynomial : P(x) = 3x^3 -2x^2 -5x-2 Give an explanation to how as well.

OpenStudy (shamil98):

@wio

OpenStudy (anonymous):

This should help: http://en.wikipedia.org/wiki/Rational_root_theorem

OpenStudy (campbell_st):

do you know if the curve actually has a rational root..?

OpenStudy (anonymous):

Basically you have to guess and check:\[ \pm \frac{1,2}{1,3} \]

OpenStudy (shamil98):

So you substitute ? for p(1), p(2)..

OpenStudy (anonymous):

So for example: \[ \frac 11, -\frac 11, \frac 21, -\frac 21, \frac 13, -\frac 13, \frac 23, -\frac 23 \]These are all potential solutions.

OpenStudy (anonymous):

So you can check each one and hope they're a root.

OpenStudy (shamil98):

That seems like a tedious job.. plugging in all those numbers especially with the equation being so long lol

OpenStudy (anonymous):

Well you know there are only 3 roots.

OpenStudy (campbell_st):

doing a quick mental arithmetic check there is a root between 1 and 2, which isn't rational

OpenStudy (anonymous):

If you find one of the roots, you can divide out \(x-r\) where \(r\) is the root. Then you have a quadratic equation.

OpenStudy (shamil98):

What does the term "rational zeros" implicate? ..

OpenStudy (campbell_st):

and doing a quick check to the left of zero... looking for change in signs, there doesn't appear to be any other roots...

OpenStudy (campbell_st):

rational means the value can be written as a fraction e.g. 0.33333... = 1/3 a rational number is any number that can be written as a/b where a and b are integers

OpenStudy (shamil98):

Okay.

OpenStudy (shamil98):

So, rational zeros are any value that is rational? lol

OpenStudy (campbell_st):

a rational zero is a value which cuts or touches the horizontal axis and is a rational number...

OpenStudy (shamil98):

thank you guys for the explanation. I'm a bit more knowledgeable of math now. :)